OpenStudy (anonymous):

For f(x)=2x + 1 and g(x)=x^2 - 7 find (f+g) (x) Answers are x^2 + 2x -6 x^2 + 2x +8 2x^3 -6 2x^2 - 15

OpenStudy (anonymous):

when you say for example: \[f(x)=4x^2+3,~~~~~and~~~~~g(x)=x^2-x+3\] then,\[\color{blue}{(f+g)(x)}=(4x^2+3)+(x^2-x+3)=4x^2+3+x^2-x+3=\color{blue}{5x^2-x+6}\]

OpenStudy (anonymous):

your problem is similar, you are adding what both of the functions are equivalent to.

OpenStudy (anonymous):

Tallan, what is not clear, ask if you have some questions.

OpenStudy (anonymous):

Im currently trying to solve it, nothing is unclear yet.

OpenStudy (anonymous):

wait ok I have a question now. Do they have to be in parantheses? the 2x+1 and everything else in order to work?

OpenStudy (anonymous):

I will set it up for you, and you will calculate: f(x)=2x + 1 g(x)=x^2 - 7 ( f + g ) (x) \\ \\ (2x+1) + (x^2-7).

OpenStudy (anonymous):

see what I am doing?

OpenStudy (anonymous):

yes! so the 7 is negative then correct?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

the x^2-7 the 7 would be a negative 7 in this equation right? Or is it a positive and were just subtracting from it?

OpenStudy (anonymous):

you are leaving out some terms. x^2-7 is just the g(x) but if you mean that you will remove the parenthesis, and from having: (2x+1) + (x^2-7). you will get: 2x+1 + x^2-7 (none of the signs change)

OpenStudy (anonymous):

ok, So when I calculated it I got A which was x^2 + 2x - 6

Nnesha (nnesha):

tallan bec this time you have to add things and its doesn't matter if you put parenthesis or not but when you have to subtract then you have to be careful like we did last time :)

OpenStudy (anonymous):

ohm, nvm you are correct it is A.

OpenStudy (anonymous):

good work Tallan!

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

not a problem:)